The indicated function y1(x) is a solution of the associated homogeneous equation. Use the method of reduction of order to find a second solution y2(x) of the homogeneous equation and a particular solution of the given nonhomogeneous equation.

y'' − 25y = 2; y1 = e^−5x

y1(x) = _____
y2(x) = ______
1
Expert's answer
2020-06-30T16:21:15-0400

Solution:

y'' − 25y = 2;

y1 = e-5x 

y'' − 25y =0

y2(x)=u(x)y1(x)

y=ue-5x

y'=u'e-5x-5ue-5x

y''=u''e-5x-10u'e-5x+25ue-5x

y'' − 25y=u''e-5x-10u'e-5x+25ue-5x-25ue-5x=0

u''e-5x-10u'e-5x=0 | e5x

u''-10u'=0

t=u'

t'-10t=0

dt/dx=10t

dt/t=10dx,

ln|t|=10x+C1

t=Ce10x

Let C=1.

y=e10xe-5x=e5x

y2(x)=e5x

yp(x)=A

A''=0

-25A=2=>A=-2/25

yp(x)=-2/25

Answer:

y2(x)=e5x

yp(x)=-2/25